Exercise. Given the equation, calcite the standard free energy change. Answer what is being asked.



1. N2(g)+3H2(g) --->2NH3(g)


If ΔH = -92.2kJ and ΔS= -0.1987kJ/K, what is ΔG for the reaction at 475°C and at 5°C? Provide the answer in kJ. In each case, is the reaction spontaneous?



2. The enthalpy and entropy change of a reaction are -3.9 kJ/mole and +56.6 J/mole K respectively at 25o C. What is the free energy change in kJ/mole? Is this reaction always spontaneous, never spontaneous or does it depend on the temperature?

1
Expert's answer
2021-11-22T17:05:03-0500

∆G = ∆H - ∆TS


∆H = -3.9KJ/mol


T = 25 + 273 = 298K


∆S = 56.6J/mol = 0.056KJ/mol


∆G = -3.9 - 298(0.056)


= -3.9 - 16.688


= -20.588KJ/mol


The reaction is spontaneous